OpenStudy (anonymous):

solve: log(x-1)+log(x-3)=8

OpenStudy (jhannybean):

Use:\[\log(a) +\log(b) = \log(ab)\]

OpenStudy (jhannybean):

\[a = (x-1) ~,~ b=(x-3)\]

OpenStudy (anonymous):

and what do i do with the 8?

OpenStudy (jhannybean):

Just worry about the left side for now. What do you gt on the left side?

OpenStudy (anonymous):



OpenStudy (jhannybean):

Good.

OpenStudy (jhannybean):

\[\log(x-1) +\log(x-3)= \log(x^2-4x+3)\] Now you have \[\log(x^2-4x+3) =8\]To get rid of the log function and solve for the variables, we can raise both sides of this function by 10.\[\large 10^{\log(x^2-4x+3)}=10^8\]

OpenStudy (jhannybean):

We are then left with : \(x^2-4x+3 = 10^8\)

OpenStudy (jhannybean):

Oh I made a big mistake.

OpenStudy (jhannybean):

Do you see what I did wrong?

OpenStudy (jhannybean):

Remember: \(\large e^{\log_e(x)} = x~ , ~\log_e(e^x) =x\)

OpenStudy (jhannybean):

So instead of raising both sides by 10, I should have actually raised both sides by base \(e\) to get rid of the log function.

OpenStudy (jhannybean):

\[\large \log(x-1) +\log(x-3)= \log(x^2-4x+3)\]\[\large \log(x^2-4x+3) = 8\]\[\large e^{\log(x^2-4x+3)} = e^8\]\[\large x^2-4x+3=e^8\]

OpenStudy (jhannybean):

Can you solve it from here, or would you like me to keep going?

OpenStudy (anonymous):

I would appreciate it, if you kept going. This concept is still confusing/new to me. Thank you so much for your help so far

OpenStudy (jhannybean):

Are you confused about any step?

OpenStudy (jhannybean):

In particular, I mean.

OpenStudy (anonymous):

Not right now. Another side question: if it was written with "ln" instead of "log", would I still use "e"?

ganeshie8 (ganeshie8):

can you take a screenshot of actual question and post it ?

ganeshie8 (ganeshie8):

in school text books, when base is omitted log generally refers to base 10 and ln always refers to base e

OpenStudy (jhannybean):

Ah, ok. I did not know how to explain it.

ganeshie8 (ganeshie8):

for school textbooks : \[\ln (x) = \log_e (x)\] \[\log(x) = \log_{10}(x)\]

OpenStudy (anonymous):

i meant. for example if the equation was: ln(x-1+ln(x-3)=8. Will I still use base e in the equation?